# Cluster Synchronization via Graph Laplacian Eigenvectors

### v1

@@ II.1.1 Eigenvectors as a Basis for Dynamics
From PDEs and solutions to the Helmholtz equation, we find intuition for understanding the dynamics on a graph through the eigenvectors of the graph Laplacian. In the solution to the heat equation or wave equation on a surface, the corresponding eigenfunctions represent ’spatial frequencies’ that are composed in a fourier series to describe solution functions on the surface [17]. The graph Laplacian can be thought of as a discretized Laplacian operator, abstracted to a graph. This spatial structure is both discrete and in general non-uniform.
The Laplacian matrix emulates diffusion across the graph, analogously to the laplace operator. The matrix consequently has a well established connection with random walks. Indeed one normalized graph Laplacian variant known as the random-walk normalized Laplacian [18]. The eigenvectors and eigenvalues can be understood similarly to the eigenfunctions in that they encode a ’spatial frequency’ but discretized over the non-uniform graph structure. One way in which we can see this reflected is by measuring zero crossings along random walks over the graph [6]. The rows and columns of the graph Laplacian are associated to the vertices of the graph. In analogy to functions, the eigenvectors of this matrix are of length $n$ and assign a value to each vertex. In considering random walks on our graph, one can count the frequency of sign changes along those walks and see positive correlation with the eigenvalue of that eigenvector [6]. This yields an intuition of ’spatial frequency’ and how it is abstracted to graphs. A relevant example is the Fiedler eigenvector, associated to the first non-zero eigenvalue of the graph Laplacian. The vertex assignement here can be thought of as a gradient from negative to positive values across the graph, which naturally presents a low ’spatial frequency’. This idea opens the door for us to construct a graph-equivalent of the fourier series solution method for PDEs, describing solutions in terms of ’spatial frequencies’ [6].
@@
Furthermore, the structural eigenvectors in the spectrum of a graph Laplacian, as shown in Corollary 2, are within the column space of $P$, assigning the same values to vertices within each partition cell. Thus the structural eigenvectors create a sub basis in the spectrum of the graph Laplacian, which allow for the total description of clustered dynamics, reducing to the quotient graph.
@@
Let $G$ be a graph, $L$ its graph Laplacian, and $\pi$ an AEP on its vertex set. The set $\{P^{\pi}v|v\text{ eigenvector of $L^{\pi}$}\}$ forms a sub-basis in the eigenvector basis of $L$ and fully describes vertex localized behavior on $\pi$.
In section II.2 we will see how the eigenvalues become relevant to the synchronization of dissipative models like the Kuramoto model. We draw from the notion of assortativity in community structure to understand the placement of structural eigenvectors in the spectrum of the Laplacian. An assortative partitioning of the vertex set has greater intra cell than inter cell connectivity. A disassortative partitioning has the converse. Moreover, the assortativity of the partitioning has been linked to the placement of the structural eigenvector in the spectrum of the Laplacian [19]. In particular, for large enough $n$, structural eigenvectors associated to assortative partitions tend towards smaller associated eigenvalues, whereas the structural eigenvectors associated to disassortative partitions tend towards the larger eigenvalues. The theory for this comes from community detection [20], but we can explain this phenomenon heuristically through the lens of AEP vertex assignments and random walks with structural eigenvectors. Given a highly assortative AEP, random walks are likely to remain inside the partition cell they start from, and since the vertex assignments of the corresponding eigenvector are constant within those cells, the random walk is likely to correspond to relatively few zero crossings in a corresponding structural eigenvector. This lower number of zero crossings for the associated structural eigenvectors suggests a lower eigenvalue. A similar argument follows for a disassortative AEP.
@@ II.2 Kuramoto Dynamics
Previous work has shown that with some persuasion, one can perform a change of variable of equation 1 to isolate the coefficients of this decomposition [21]. We extend this work to include weighted networks and antisymmetric phase offsets. In this section, we only consider the weighted extension, seen below, and we study the latter in section IV.2.
@@
$\alpha_{0}$ $=\sqrt{N}\overline{\omega}t$ (3)
@@
We call this basis change of the Kuramoto equations, the coefficient equations. This transformation frames the problem of synchronization in terms of the population dynamics of coefficients. When the system reaches a globally locked frequency, then naturally the zero mode describes the advancement of all phase angles, while all other coefficients describe the phase offsets of the system. Thus in this form of synchronized behavior, the coefficients for non-zero eigenvalues approach constant values [21]. Assuming coefficients are small and close to synchronization, we can solve a linearization of equation 2. For $r\neq 0$:
@@
We note that the form of this equation does not take into consideration the weights of the graph. All structural information related to edge weights, impacting the synchronization in the system, is encoded in the eigenvalues and eigenvectors of the graph Laplacian. Thus this limit is significant in that it reflects the structure present in the system.
@@ II.3 Spectral Kuramoto Clustering
Previous work has shown that a weighted AEP can induce cluster synchronized behavior in the Kuramoto model [22, 23, 4]. We use the spectral properties of AEPs to recast this behavior in the context of the coefficient equation. This framework provides us with inroads to explore more general synchronized behavior.
Using the coefficient end behavior in equation 5, we can describe the impact of partitions and graph structure on synchronization. If the natural frequency vector has constant values within the cells of an AEP, then by corollary 4, $\vec{\omega}$ is described completely within the scaled up eigenbasis of the corresponding quotient Laplacian. By orthogonality, all eigenvectors not from the quotient Laplacian must be orthogonal to $\vec{\omega}$. Thus the linearized coefficient limit is only nonzero for the structural eigenvectors. As a result, only the structural eigenvectors are nonzero in the end-state eigendecomposition of the system (fig 2). Expressed in terms of the dynamics of the original Kuramoto system, this suggests that the system experiences cluster synchronization, where the phases of oscillators within the clusters of the AEP remain equal. This aligns with work such as [24] which has found that synchronized clusters in similar models of chaotic oscillators require identical elements in some eigenvector (fig 2).
We note that since equation 5 is from a linearization of the system, the approximation it gives diverges for large end state values of coefficients. For fully synchronized systems, the range of error is proportional to the coefficient in question (fig 3). Thus the modes orthogonal to the natural frequency vector respect the approximation, and the qualitative dynamics as a whole remain the same.
Figure 2: This example highlights the impact of structural eigenvectors on clustered synchronization. A. depicts a $15$ vertex graph, along with a color coded AEP which it admits. We place a Kuramoto model on this graph with uniform natural frequencies within the AEPs cells. The resulting phase plots in B. show a synchronization to a full phase clustered regime. The corresponding eigendecomposition is shown in C.. Cluster synchronization corresponds to the dominance of the structural sub-basis, in this case a single structural mode corresponding to $\lambda_{1}$.
@@
Figure 3: Here we see exemplified that the linearized approximations of the synchronized end state coefficients diverge in proportion to the coefficients’ magnitude.
@@ II.4 Transient Cluster Synchronization
$\vec{\theta}(t)=\alpha_{0}(t)v_{0}=\sqrt{N}\delta t\mathbb{1}$
@@
Using the equations 4, we can see the path the eigenmodes take to equilibrium. We observe that in the described setting, for $r\neq 0$ the solution follows $\alpha_{r}(t)=\alpha_{r}(0)e^{-\sigma\lambda_{r}t}$. This indicates faster convergence for the larger eigenvalue modes. In other words, with comparable initial values, a difference in eigenvalues leads to the separation of convergence rates to synchronization. This explains the observations of works such as [25] or [8], which found convergence to synchrony as being in clusters, with delays in transient clustering regimes being proportional to the gaps in the eigenspectrum of the Laplacian. If the network has non trivial AEPs, those would show up in the eigenspectrum and if their placement in the graph Laplacian spectrum is low, then their coefficients will be the slowest to approach zero (fig 4 C.). In the presence of an AEP, we define cluster synchronization as a restriction of the dynamics to the structural sub-basis. Using corollary 4 this means that the described scenario results in transient clustered behavior while approaching equilibrium. The key here is that gaps in the eigenspectrum can relate to regime changes when approaching synchronization because the rate of asymptotic decay to the limit of a coefficient is determined by the eigenvalue of that mode. IT is summarized in proposition 5.
@@
Let $G$ be a graph with an AEP, $\pi$. Suppose $\pi$ is sufficiently assortative in that the $p\geq 1$ smallest eigenvalues of the graph Laplacian, $L$, are associated to structural eigenvectors. Assume there are no repeated eigenvalues. Consider a Kuramoto model on $G$ with uniform natural frequencies. Then up to the choice of initial condition, in the linearized solution the system will exhibit transient cluster synchronization along $\pi$.
Further, since AEPs can occur at multiple scales, this can also describe transitions through multiple clustering regimes. In a graph with assortative hierarchical clustering, the structural eigenvectors of coarser AEPs generically have lower eigenvalue than the finer ones. If assortative enough, these structural eigenvectors occupy the lower frequency modes. While synchronizing this would lead to higher frequency, non structural, modes dying initially to reveal the more slowly decaying structural modes, whose overlap would manifest as a finely clustered behavior. Then the lower frequency structural modes dominate in time as well. The duration of these separate clustering regimes is proportional to the gap in the eigenspectrum between the sequential modes. The following result naturally follows from the previous.
@@
Let $G$ be a graph with assortative $k-$layer hierarchical AEP clustering; that is, nested AEPs occupying the $p\geq k$ smallest eigenvalues of the graph Laplacian spectrum in ascending order of coarseness. Assume there are no repeated eigenvalues. Consider a Kuramoto model on $G$ with uniform natural frequencies. Then up to the choice of initial condition, in the linearized solution the system will exhibit multiple regimes of transient cluster synchronization.
Figure 4: A. A hierarchically clustered network of $180$ vertices.B. A phase plot showing the convergence to total synchronization. We observe $4$ distinct states. Initial disordered behavior (time steps 0-50), followed by clustering into $6$ clusters (time steps 50-350). After this, the $6$ clusters merge into $3$, which slowly decay toward merging into a single fully synchronized component. C. We see these $4$ states represented in the eigendecomposition. An initial spread of coefficients representing disordered behavior, before the nonstructural ones ($\alpha_{r}$, $r>5$) decay, leaving only structural ones. Then the higher eigenvalue structural modes decay leaving just those corresponding the coarsest AEP.

### v2

@@ II.1.1 Eigenvectors as a Basis for Dynamics
From PDEs and solutions to the Helmholtz equation, we find intuition for understanding the dynamics on a graph through the eigenvectors of the graph Laplacian. In the solution to the heat equation or wave equation on a surface, the corresponding eigenfunctions represent ’spatial frequencies’ that are composed in a Fourier series to describe solution functions on the surface.Haberman (2013) The graph Laplacian can be thought of as a discretized Laplacian operator, abstracted to a graph. This spatial structure is both discrete and in general non-uniform.
The Laplacian matrix emulates diffusion across the graph, analogously to the Laplace operator.Belkin and Niyogi (2008) The matrix has (together with its deformed variants) a well established connection with the diffusive process of random walks on graphs.Seabrook and Wiskott (2023); Boukrab and Pagès-Zamora (2021) Further, its eigenvectors and eigenvalues encode ’spatial frequency’, playing a similar role to the eigenfunctions of the Helmholtz equation in the case of diffusion on a surface, but discretized over a non-uniform graph structure. One way in which we can see this reflected is by measuring zero crossings along random walks over the graph.Shuman et al. (2013) The rows and columns of the graph Laplacian are associated to the vertices of the graph. In analogy to functions, the eigenvectors of this matrix are of length $n$ and assign a value to each vertex. In considering random walks on our graph, one can count the frequency of sign changes along those walks and see positive correlation with the eigenvalue of that eigenvector.Shuman et al. (2013) This yields an intuition of ’spatial frequency’ and how it is abstracted to graphs. A relevant example is the Fiedler eigenvector, associated to the first non-zero eigenvalue of the graph Laplacian. The vertex assignement here can be thought of as a gradient from negative to positive values across the graph, which naturally presents a low ’spatial frequency’. This idea opens the door for us to construct a graph-equivalent of the Fourier series solution method for PDEs, describing solutions in terms of ’spatial frequencies’.Shuman et al. (2013)
@@
Furthermore, the structural eigenvectors in the spectrum of a graph Laplacian, as shown in Corollary 2, are within the column space of $P$, and therefore assign the same values to vertices within each partition cell. Thus the structural eigenvectors define a sub-basis in the spectrum of the graph Laplacian that captures all information relevant to the clustered dynamics, allowing the system to be effectively described by the quotient graph. We can formalize this conclusion in the following statement.
@@
Let $G$ be a graph, $L$ its graph Laplacian, and $\pi$ an AEP on its vertex set. The set $\{P^{\pi}v|v\text{ eigenvector of $L^{\pi}$}\}$ forms a sub-basis of the eigenvector basis of $L$ which spans the space of vectors that are constant on the cells of $\pi$, and therefore encodes all dynamics constrained by the partition.
Proof.
Given the graph $G$ and an AEP $\pi$ of size $k$ on its vertex set, corollary 2 guarantees that $B=\{P^{\pi}v|v\text{ eigenvector of $L^{\pi}$}\}$ is a subset of eigenvectors of the graph Laplacian. By proposition 3, the set of eigenvectors of $L^{\pi}$ span $\mathbb{R}^{k}$. It follows that $B$ spans the column space of $P^{\pi}$. Recall that vectors in this column space are of size $n$ and have constant values for all entries corresponding to vertices within the same cell of $\pi$. Thus $B$ is a subspace of the eigenspace of $L$ which spans all vector valued vertex assignments constant within the cells of $\pi$ in $G$. ∎
In section II.2 we will see how the eigenvalues become relevant to the synchronization of dissipative models like the Kuramoto model. We draw from the notion of assortativity in community structure to understand the placement of structural eigenvectors in the spectrum of the Laplacian. An assortative partitioning of the vertex set has greater intra cell than inter cell connectivity. A disassortative partitioning has the converse. Moreover, the assortativity of the partitioning has been linked to the placement of the structural eigenvector in the spectrum of the Laplacian.Krzakala et al. (2013) In particular, for large enough $n$, structural eigenvectors associated to assortative partitions tend towards smaller associated eigenvalues, whereas the structural eigenvectors associated to disassortative partitions tend towards the larger eigenvalues. The theory for this comes from community detection,Saade, Krzakala, and Zdeborová (2014) but we can explain this phenomenon heuristically through the lens of AEP vertex assignments and random walks with structural eigenvectors. Given a highly assortative AEP, random walks are likely to remain inside the partition cell they start from, and since the vertex assignments of the corresponding eigenvector are constant within those cells, the random walk is likely to correspond to relatively few zero crossings in a corresponding structural eigenvector. This lower number of zero crossings for the associated structural eigenvectors suggests a lower eigenvalue. A similar argument follows for a disassortative AEP.
@@ II.2 Kuramoto Dynamics
Previous work has shown that with some persuasion, one can perform a change of variable of equation 1 to isolate the coefficients of this decomposition.Kalloniatis (2010) We extend this work to include weighted networks and antisymmetric phase offsets. In this section, we only consider the weighted extension, seen below, and we study the latter in section IV.2,
@@
$\alpha_{0}$ $=\sqrt{N}\overline{\omega}t.$ (3)
@@
We call this basis change of the Kuramoto equations, the coefficient equations. This transformation frames the problem of synchronization in terms of the population dynamics of coefficients. When the system reaches a globally locked frequency, then naturally the zeroth mode describes the advancement of all phase angles, while all other coefficients describe the phase offsets of the system. Thus in this form of synchronized behavior, the coefficients for non-zero eigenvalues approach constant values.Kalloniatis (2010) Assuming coefficients are small and close to synchronization, we can solve a linearization of equation 2. For $r\neq 0$:
@@
We note that the form of this equation does not require the weights of the graph. All structural information related to edge weights, impacting the synchronization in the system, is encoded in the eigenvalues and eigenvectors of the graph Laplacian. Thus this limit is significant in that it reflects the structure present in the system.
@@ II.3 Spectral Kuramoto Clustering
Previous work has shown that a weighted AEP can induce cluster synchronized behavior in the Kuramoto model.Kato and Ishii (2023); Menara et al. (2020); Schaub et al. (2016) We use the spectral properties of AEPs to recast this behavior in the context of the coefficient equation. This framework provides us with inroads to explore more general synchronized behavior.
Using the coefficient end behavior in equation 5, we can describe the impact of partitions and graph structure on synchronization. If the natural frequency vector has constant values within the cells of an AEP, then by corollary 4, $\vec{\omega}$ is described completely within the scaled up eigenbasis of the corresponding quotient Laplacian. By orthogonality, all eigenvectors not from the quotient Laplacian must be orthogonal to $\vec{\omega}$. Thus the linearized coefficient limit is only nonzero for the structural eigenvectors. As a result, only the structural eigenvectors are nonzero in the end-state eigendecomposition of the system (figure 2). Expressed in terms of the dynamics of the original Kuramoto system, this suggests that the system experiences cluster synchronization, where the phases of oscillators within the clusters of the AEP remain equal. The significance of this choice is that the natural frequency distribution over the vertices excites the structural modes rather than the nonstructural ones, allowing for these exact clustered dynamics. This aligns with work such asFan and Wang (2023) which has found that synchronized clusters in similar models of chaotic oscillators require identical elements in some eigenvector (figure 2).
We note that since equation 5 is from a linearization of the system, the approximation it gives diverges for large end state values of coefficients. For fully synchronized systems, the range of error is proportional to the coefficient in question (figure 3). Thus the modes orthogonal to the natural frequency vector respect the approximation, and the qualitative dynamics as a whole remain the same.
Figure 2: This example highlights the impact of structural eigenvectors on clustered synchronization. A. The network consists of 15 vertices, color-coded according to an almost equitable partition (AEP) of the network. A Kuramoto model is placed on the graph, with coupling $\sigma=2$ and natural frequencies set to be identical within each cell of the AEP and distinct across cells (respectively the pink, yellow, blue cells have frequencies 3,4,5). This frequency assignment reinforces the structural partition. B. The resulting phase dynamics show convergence to a phase-clustered regime consistent with the AEP. C. Time evolution of the projection coefficients onto the Laplacian eigenbasis. Cluster synchronization corresponds to the dominance of the structural subspace, here captured primarily by $\alpha_{1}$, the eigenmode associated with $\lambda_{1}$, whose asymptotic limit is approximated by $\alpha_{1}^{\infty}$.
@@
Figure 3: An example to highlight the convergence behavior of spectral coefficients using the toy network presented in figure 2.A with the same AEP on the vertices but different parameters in the Kuramoto model. In this example, we set the natural frequencies to be identical within each cell of the AEP and distinct across cells (respectively the pink, yellow, blue cells have frequencies 3,4,9), and we increased the coupling strength between oscillators to 5. which leads to the coeffients $\alpha$ having a slightly different behavior than the ones in figure 2.C. We see that the deviation between the time-dependent coefficients and their linearized limit values decreases with the magnitude of the corresponding mode. In particular, $\alpha_{1}(t)$ converges more closely to $\alpha_{1}^{\infty}$ than the higher-index mode $\alpha_{2}(t)$, exemplifying this behavior.
@@ II.4 Transient Cluster Synchronization
$\vec{\theta}(t)=\alpha_{0}(t)v_{0}=\sqrt{N}\delta t\mathbb{1}.$
@@
Using the equation 4, we can see the path the eigenmodes take to equilibrium. We observe that in the described setting, for $r\neq 0$ the solution follows $\alpha_{r}(t)=\alpha_{r}(0)e^{-\sigma\lambda_{r}t}$. This indicates faster convergence for the larger eigenvalue modes. In other words, with comparable initial values, a difference in eigenvalues leads to the separation of convergence rates to synchronization. This explains the observations of previous works,De Domenico (2017); Arenas, Díaz-Guilera, and Pérez-Vicente (2006) which found convergence to synchrony as being in clusters, with delays in transient clustering regimes being proportional to the gaps in the eigenspectrum of the Laplacian. If the network has non trivial AEPs, those would show up in the eigenspectrum and if their placement in the graph Laplacian spectrum is low, then their coefficients will be the slowest to approach zero (figure 4 C.). In the presence of an AEP, we define cluster synchronization as a restriction of the dynamics to the structural sub-basis. Using corollary 4 this means that the described scenario results in transient clustered behavior while approaching equilibrium (figure 4 B). The key here is that gaps in the eigenspectrum can relate to regime changes when approaching synchronization because the rate of asymptotic decay to the limit of a coefficient is determined by the eigenvalue of that mode. It is summarized in proposition 5, with the proof in appendix B.
@@
Let $G$ be a connected graph with an AEP, $\pi$. Suppose $\pi$ is sufficiently assortative in that the $p\geq 1$ smallest eigenvalues of the graph Laplacian, $L$, are associated to corresponding structural eigenvectors. Assume there are no repeated eigenvalues. Consider a Kuramoto model on $G$ with uniform natural frequencies, and assume $\alpha_{i}(0)>0$ for some $0<i\leq p$, Then in the linearized solution the system will exhibit transient cluster synchronization along $\pi$.
As usual the above clustered behavior is defined as the domination of the structural sub-basis. However, it is important to note the impact of the choice of initial condition. All initial conditions satisfying 5 will eventually be dominated by the assortative structural eigenmodes, but choices in which the structural modes have small initial condition will result in all nonzero modes being small by the time the structural modes dominate. In the phase dynamics this is reflected as clustered dynamics where the clusters are very close in phase to one another.
Since AEPs can occur at multiple scales, this can also describe transitions through multiple clustering regimes. Consider a graph $G$ with a sequence of possible almost equitable partitions $\pi_{1},\pi_{2},\dots,\pi_{k}$ such that $\pi_{1}\prec\pi_{2}\prec\dots\prec\pi_{k}$, meaning each partition is a refinement of the next (i.e., each cell in $\pi_{i}$ is fully contained in a cell in $\pi_{i+1}$). We say this sequence defines a $k$-layer hierarchical AEP clustering of $G$.
In a graph with assortative hierarchical clustering, i.e. each $\pi_{i}$ is assortative as defined in section II.1.1, then the structural eigenvectors of coarser AEPs generically have lower eigenvalue than the finer ones.Krzakala et al. (2013); Schaub, Li, and Peel (2023) If assortative enough, these structural eigenvectors occupy the lower frequency modes (see an example of such a network in figure 4 A). While approaching full synchronization this would lead to higher frequency, non structural, modes dying initially to reveal the more slowly decaying structural modes, whose overlap would manifest as a finely clustered behavior (see lemma 12 in appendix B). Then in time the lower frequency structural modes dominate the higher frequency structural modes as well. Assuming similar starting values, the duration of these separate clustering regimes is proportional to the gap in the eigenspectrum between sequential modes. The following result naturally follows from the previous, with the full proof in appendix B and example in figure 4.
@@
Let $G$ be a graph with $k$-layer hierarchical AEP clustering defined by a sequence of nested partitions $\pi_{1}\prec\pi_{2}\prec\dots\prec\pi_{k}$. Assume assortativity such that for some choice of indices $0=p_{0}<p_{1}<\cdots<p_{k}<n$, for each $0<i\leq k$ the eigenvalues $\lambda_{p_{i-1}+1}<\dotsc<\lambda_{p_{i}}$ correspond to structural eigenvectors of $\pi_{i}$. Assume all eigenvalues are simple (non-repeated). Consider a Kuramoto model on $G$ with uniform natural frequencies. Then up to the choice of initial condition, in the linearized regime the system will pass through a sequence of transient cluster synchronization states corresponding to each level of the AEP hierarchy, before ultimately converging to global synchrony.
Figure 4: Hierarchically related structural modes. A. A hierarchically clustered network with 180 vertices. Vertex positions illustrate one of the allowable AEPs with 6 clusters, while vertex and edge coloring highlight a coarser AEP consisting of 3 clusters. To ensure full system synchronization, the natural frequencies were uniformly set to 3 across all vertices along with a coupling of $\sigma=0.3$. B. Phase dynamics showing convergence to full synchronization. Four distinct dynamical stages are observed: an initial disordered regime (time steps 0–50), followed by the formation of 6 clusters (time steps 50–350); these then merge into 3 larger clusters, which gradually converge into a single synchronized group. C. Time evolution of projection coefficients onto the Laplacian eigenbasis, illustrating the same sequence of transitions. Initially, a wide distribution of coefficients reflects disorder; this is followed by the decay of nonstructural modes ($\alpha_{r}$, $r>5$), leaving only the structural components ($\alpha_{r}$, $r\leq 5$). Subsequently, the finer-scale structural modes decay, and the dynamics are eventually dominated by an eigenmode corresponding to the coarsest AEP, $\alpha_{1}$. The ultimate convergence to global phase synchronization is reflected by all coefficients asymptotically approaching their limits $\alpha_{r}^{\infty}=0$ for all $r$, a result enabled by the uniform assignment of natural frequencies across all vertices.